If bacteria are transformed with 1ug of DNA and 76 CFU result from spreading 100ul of the culture (from 500ul final volume), what is the transformation efficiency (in CFU per ug of DNA)?
Answer #1

Transformation efficiency = number of colonies transformed / amount of DNA spread on plate

Amount of DNA spread on plate = volume spread × amount of DNA / total volume of transformation = 100 × 1 / 500 = 0.2

Number of colonies = 76

Transformation efficiency = 76 / 0.2 = 380 CFU per microgram

Please rate high.